OpenStudy (anonymous):

Finish writing the equation of the line y=-1/3x- ____, given the points (3, -5) and (-6, -2). Type only the part that is missing in the equation

OpenStudy (anonymous):

y=-1/3x- ____, y=mx+(-c) -2=1/3(-6)+(-c) -2+2=c 0=c y=-1/3x- ____, -5=-1/3(3)-c -5+1=-c -4=-c 4=c